OpenStudy (anonymous):

What is the solution to the following system of equations? y = x^2 + 5x + 3 y = x^2 + 3x – 3 A. (–3, 3) B. (5, 3) C. (5, –3) D. (–3, –3)

OpenStudy (anonymous):

(-3 , -3)

OpenStudy (anonymous):

x^2 + 5x + 3 = x^2 + 3x - 3 5x+3 = 3x -3 2x = -6 ---> x = -3 ---> y = (-3)^2 + 5*(-3) +3 = -3